OpenStudy (11study11):

A student is training for the swim team during July and August one summer. She swims one hour every day, while averaging 800 meters each day. Select the graph that correctly shows the total meters swam during the time of her training program. Graph of meters versus hours. Ray starting at point zero comma zero and passing through 24000 comma 30. Graph of meters versus hours. Line segment between point zero comma zero and 49 comma 62. Graph of hours versus meters. Line segment between point zero comma zero and 62 comma 49,600. Graph of hours versus meters. Ray starting at point zero c

OpenStudy (11study11):

@jim_thompson5910

OpenStudy (11study11):

is it D?

OpenStudy (11study11):

@Data_Lg2

OpenStudy (-studywith-):

First you need to decide which is x and y: hours or distance? Well x is the independent and y is the opposite so... The distance depends on the number of hours they swam, right? So the distance is y and the hours (time) is x.

OpenStudy (-studywith-):

Then you can decide the points, right?

OpenStudy (-studywith-):

Next you need to decide the number of hours she swam for the x-coordinate. Well, she swam an hour every day for these two months, how many days are between July and August?

OpenStudy (11study11):

like 30

OpenStudy (-studywith-):

2 months? That's more like 1 month.

OpenStudy (11study11):

oh 60

OpenStudy (-studywith-):

Remember, July and August both have 31 days. 31 x 2 = 62.

OpenStudy (11study11):

yup

OpenStudy (11study11):

so it is D?

OpenStudy (-studywith-):

Yeah, since 62 is the hours (time) which is the x-coordinate. Don't get confused with B, that's 62 as the y-coordinate which should be distance. But yeah, you're lucky you don't have to do any more work than that.

OpenStudy (11study11):

Thank you

OpenStudy (11study11):

can you check one more ?

OpenStudy (-studywith-):

Yeah, of course. Thanks for the medal, btw

OpenStudy (11study11):

A student is running a 5-kilometer race. He runs 1 kilometer every 3 minutes. Select the function that describes his distance from the finish line after x minutes. f(x) = − 1 over 3 x + 5 f(x) = − 1 over 5 x + 3 f(x) = 1 over 5 x + 5 f(x) = 1 over 3 x + 3

OpenStudy (-studywith-):

So x is the mins, right?

OpenStudy (11study11):

yup

OpenStudy (-studywith-):

So y is the distance, aka km?

OpenStudy (11study11):

ok

OpenStudy (-studywith-):

Wait, no f(x) is, one sec I haven't done these in a while sorry.

OpenStudy (-studywith-):

x is the function of y so f(x) is y

OpenStudy (-studywith-):

You still with me?

OpenStudy (11study11):

ohh

OpenStudy (-studywith-):

Ok, so you know y = mx + b?

OpenStudy (-studywith-):

Well f(x) is basically y so as function notation it's the same thing as f(x) = mx + b, ok?

OpenStudy (11study11):

ok

OpenStudy (-studywith-):

Wait, you have to get it, do you? Because I can explain. Slope-intercept form is y = mx + b where m is the slope and b is the y-intercept.

OpenStudy (11study11):

yes (F)x is the same as y

OpenStudy (11study11):

f(x)

OpenStudy (-studywith-):

you get the equation, though?

OpenStudy (-studywith-):

ok, anyway find your xs and ys, then you can find m and b which is the goal.

OpenStudy (11study11):

so m is the slope and b is the y intercept right.

OpenStudy (11study11):

@-StudyWith-

OpenStudy (11study11):

is it a or c?

OpenStudy (-studywith-):

one sec

OpenStudy (-studywith-):

Let's find m first. The slope. To do that, we have to have 2 points. Well we already have one: (3, 1) If you add 2 to each coordinate it is equal so... (5, 3). There, two points.

OpenStudy (-studywith-):

You know the slope formula? y2 - y- / x2 - x1

OpenStudy (-studywith-):

Pick 2 ys and 2 xs but keep which one is 1 and 2 consistent.

OpenStudy (-studywith-):

Listen I gotta go but I'll explain it real quick.

OpenStudy (-studywith-):

We have (3, 1) and (5, 3). y1 = 1 and y2= 3. x1 = 3 and x2 = 5. 3-1/5-3= 2/2= 1 so our m is 1. so far we have f(x) = (1)x + b

OpenStudy (-studywith-):

you pick 2 random coordinates for each y and x

OpenStudy (11study11):

1,1 3,3

OpenStudy (-studywith-):

But it's from the finish line. So 1/5, right, 1 out of 5km in total?

OpenStudy (-studywith-):

Anyway, the correct answer is: f(x)=5−1/3x

OpenStudy (11study11):

so a

OpenStudy (-studywith-):

Because it's 1 km per 3 mins, and yes. I gtg sorry can't explain more.

OpenStudy (11study11):

thanks

OpenStudy (-studywith-):

It's negative because it's getting closer to the finish line each minute so distance is being taken away, make sense? And np.

OpenStudy (-studywith-):

Nevermind!